In 1998, the total demand for U.S. wheat was Q = 3244 − 283P and the domestic supply was QS = 1944 + 207P. At the end of 1998, both Brazil and Indonesia opened their wheat markets to U.S. farmers. Suppose that these new markets add 200 million bushels to U.S. wheat demand. What will be the free-market price of wheat and what quantity will be produced and sold by U.S. farmers?
1
Expert's answer
2020-12-14T05:40:53-0500

If Brazil and Indonesia add an additional 200 million bushels of wheat to U.S. wheat demand, the new demand curve ′Q"_{d}" , would be equal to Q"_{d}" + 200, or′Q"_{d}" = (3244 - 283P) + 200 = 3444 - 283P.


Equating supply and the new demand, we may determine the new equilibrium price, 1944 + 207P = 3444 - 283P, or 490P = 1500, or P* = $3.06 per bushel.

To find the equilibrium quantity, substitute the price into either the supply or demand equation, Q"_{s}" = 1944 + (207)(3.06) = 2,577.67 and Q"_{d}" = 3444 - (283)(3.06) = 2,577.67